Determine fixed, variable, and mixed costs. E56 PCB Corporation manufactures a single product. Monthly production costs incurred in the manufacturing process are shown below for the production of 3,000 units. The util ities and maintenance costs are mixed costs. The fixed portions of these costs are $300 and $200, respectively. (LO 1), AP Production in Units Production Costs Direct materials Direct labor Utilities Property taxes Indirect labor Supervisory salaries Maintenance Depreciation 3,000 7,500 18,000 2,100 1,000 4,500 1,900 1,100 2,400 Instructions (a) Identify the above costs as variable, fixed, or mixed.
Answer #1

Requirement 'a'

--Variable Costs are those cost that change in totality when the production level changes. However they remain same 'per unit'.

--Fixed Costs are those cost that are fixed and do not change, whether production volume increases or decreases.

--Mixed Cost is that cost that has the features of both - variable as well as fixed cost. This means, part of it will be fixed, and rest of it will behave in variable cost fashion.

Direct Materials Variable Cost [7500/3000 units = $2.5 per unit]
Direct Labor Variable Cost [1800/3000 units = $6 per unit]
Utilities Mixed Cost
Property Taxes Fixed Cost
Indirect Labor Variable Cost [4500/3000 units = $1.5 per unit]
Supervisory Salaries Fixed Cost
Maintenance Mixed Cost
Depreciation Fixed Cost